Here is a mixture problem. I must have a mental block against work problems because I never under stand how to set them up. I know mixture problems have to same formula to work them out, but I keep getting the numbers in the wrong places. Thanks for any help. Any suggestions would be great!!!
Mean Melinda's radiator only has 1.5L in it and it is only 33 1/3% anti-freeze. The manufacturer recommends that the radiator be filled with 4L of 50% antifreeze. How much antifreeze and how much water must Melinda add to her radiator to meet the manufacturer's specifications?
I say just take it to Jiffy lube!!!
Can you please help me?
Donna

Answer by longjonsilver(2297)

Let x = volume of antifreeze to add.

So, BEFORE = AFTER

1.5*0.333 + x*1.000 = (x+1.5)*0.50
0.5 + x = 0.50x + 0.75
0.50x = 0.25
x = 0.5 litres

So, to make up a 50% concentration requires 0.5 litres of antifreeze.
So now we have a radiator with 2 litres liquid in it.

Recommendation is for 4 litres, so we need a further 2 litres of liquid. To keep this at 50%water-50% antifreeze, we add 1 litre of each.

So, in total we have added 1.5 litres of antifreeze and 1 litre of water.

Does this make sense?

jon.